OpenStudy (anonymous):

how to intargrate e^-x^2

OpenStudy (abb0t):

you can expand the exponential using power series to get a function.

OpenStudy (anonymous):

Or you can use get it into another dimension as well. \[\iint_R e^\left\{ -x^2-y^2 \right\}d\]

OpenStudy (abb0t):

@Spacelimbus is right, you can also do that! Convert it to polar coordinates, I think.

OpenStudy (anonymous):

sorry, the equation editor messed up for me, it should be dydx

OpenStudy (anonymous):

\[\iint_R e ^\left\{ -x^2-y^2 \right\}dxdy\]

OpenStudy (anonymous):

And from there, as @abb0t mentioned, get it into polar coordinates.

OpenStudy (anonymous):

thanks guys i already tryied by parts

OpenStudy (anonymous):

\[\iint_R r e^ \left\{ -r^2 \right\}drd \theta\]

OpenStudy (anonymous):

Do you know how to go from here @Rhandzu ?

OpenStudy (anonymous):

@Spacelimbus , no can you please help me step-by-step

OpenStudy (abb0t):

convert to polar from where spacelimbus pointed out up there \(r^2 = x^2+y^2\) and don'4 forget about the jacobian

OpenStudy (abb0t):

You should get: \[\int\limits_{0}^{2\pi}\int\limits_{0}^{\infty} re^{-r^2}drd \theta\]

OpenStudy (abb0t):

NOW you can use u-substitution.

OpenStudy (anonymous):

\[\checkmark\]

OpenStudy (abb0t):

Which i am sure you can easily finish the problem from here.

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

i'm new to culculas but i do have math basics

OpenStudy (anonymous):

@Spacelimbus , can you please tell me the name of that formula

OpenStudy (anonymous):

Hmm good question actually, it's a technique I have once seen, I guess it's a "Double improper integral method", Dirichlet Integrals use similar methods to be evaluated. Let me see, this might help you as well: http://www.youtube.com/watch?v=fWOGfzC3IeY As seen on MIT in their Multivariable Calculus Courses.

OpenStudy (anonymous):

wow thanks great Video, so do i use the same method to integrate sin(x^2)

OpenStudy (anonymous):

no, for sin(x^2) the easiest would be to use a maclaurin series representative, the method worked above because of the exponential laws, with the trigonometric functions that wont work so easily.

OpenStudy (abb0t):

I would approach that one by using taylor series, but I think you can probably do the same.

OpenStudy (anonymous):

@abb0t tell me more

OpenStudy (anonymous):

g2g thank you very much u are great to work with,please visit my blog learntoboubleyourmoney.blogspot.com pls don'nt mind the spelling, just to say thanks my email is rrmdhluli@gmail.com